OpenStudy (linn99123):

Leo can order a small, medium, or large yogurt. There are 10 flavors and 5 toppings from which to choose. How many ways can Leo order a yogurt? A. 18 B. 35 C. 50 D. 150

OpenStudy (anonymous):

I think its C.

OpenStudy (anonymous):

i think its D

OpenStudy (cwrw238):

yes its D for each size there 5*01 = 50 combinations as there are 3 sizes we have 3*50 = 150

OpenStudy (anonymous):

I think this because you can give each of the ten flavors 5 different topping and that equals out to 50 already, then you can give each of those flavors with the different toppings 3 different sizes 3 times.

OpenStudy (cwrw238):

* 5 *10 = 50